# Synaptic-Like Plasticity in 2D Nanofluidic Memristor from Competitive Bicationic Transport

### v1

@@ Abstract
We have demonstrated synaptic-like plasticity of aqueous ion transport through a sub-nanoporous 2D membrane made of a monolayer hBN featuring an array of B3N vacancies in a binary aquous salt. The ion conductance of this nanofluidic memristor can be dynamically modulated by few-nanoseconds-long transmembrane voltage pulses, resulting in a switching speed of approximately 0.1 GHz. The resulting energy dissipation per voltage spike is very low, roughly 0.1—100 aJ per spike, which is several orders of magnitude lower than that of the human brain, 0.1—10 fJ per spike. Notably, this synaptic-like plasticity arises from the dynamic interplay between two cationic species (Na+ and K+), exemplifying an illustrative approach to leveraging ionic diversity for achieving neuromorphic functionality. The conductive state of the membrane is shown to be governed by the adsorption/desorption dynamics of Na+ trapped in the pores, while K+ ions permeate through the empty pores. We derived a set of illustrative analytical expressions associated with the apparent synaptic-like potentiation and memory volatility. Further experimentation is needed to demonstrate the predicted behaviors for various subnanoporous 2D membranes, as well as ionic permeants. It is our hope to stimulate further work aimed at a better understanding of the effects of ionic diversity, including at the system level, i.e., in the case of interconnected nanofluidic memristive elements.
I Methods
@@
The simulated system consisted of a subnanoporous hBN monolayer suspended by its perimeter inside a rectangular box filled with atomistically resolved aqueous salt mixture. The pore array consisted of 16 triangular B3N vacancies within a 6.998 nm by 6.926 nm sheet. Inside the box, the membrane plane was located at $z=h/2$, where $h=10$ nm is the box height along the $Z$-direction. The edge atoms of the membrane were tethered to their initial positions by spring restraints. The membrane was simulated using parameters developed earlier Govind Rajan et al. (2018) within the OPLS-AA forcefield framework Jorgensen et al. (1996). The partial charges of the nitrogen atoms at the edges of B3N vacancies were set at $2/3$ of their bulk values obtained earlier from quantum-chemical calculations Govind Rajan et al. (2018), ensuring electrical neutrality of the membrane. Water was simulated using the TIP4P model Jorgensen et al. (1983). All of the non-bonded interactions were simulated using the OPLS-AA forcefield framework Jorgensen et al. (1996). A particle-particle–particle-mesh scheme was used to simulate the electrostatic interactions. A 1.2 nm cut-off radius was used for all short-range interactions, including short-range Coulomb interactions and the Lennard-Jones interactions. To initialize the conductive weight of the membrane to zero in all of our simulations, we placed Na+ ions inside the B3N vacancies. The system first underwent static energy minimization and then dynamics relaxation in the NPT ensemble at $T=300$ K and $P=1$ bar, with the barostat modifying simulation box dimensions only in the $Z$-direction. Relaxed systems underwent ion transport simulations under rectangular pulses of external electric field applied in the $Z$-direction. The corresponding pulse magnitudes were calculated as the electric field magnitude, multiplied by the box height along $h$. The ion current and PMF data were obtained using procedures identical to those in our previous work Noh and Smolyanitsky (2024a). All of the MD simulations were performed using GPU-accelerated GROMACS Abraham et al. (2015); Páll et al. (2020) and the molecular visualization tasks were carried using the OVITO software Stukowski (2009).
II Acknowledgments
@@
This work was supported in part by the research grant awarded by the National Science Foundation (NSF) under Award No. 2110924, and by computational resources from NSF-supported ACCESS program under Award No. NNT230006. We acknowledge that a substantial portion of the simulations was conducted using the GPU-accelerated high-performance computing facilities at the National Institute of Standards and Technology.

### v2

@@ Abstract
To summarize, we have demonstrated synaptic-like plasticity of aqueous ion transport through a sub-nanoporous 2D membrane made of a monolayer hBN featuring an array of B3N vacancies in a binary aquous salt. The ion conductance of this nanofluidic memristor can be dynamically modulated by few-nanoseconds-long transmembrane voltage pulses, resulting in a switching speed of approximately 0.1 GHz. The resulting energy dissipation per voltage spike is very low, roughly 0.1—100 aJ per spike, which is several orders of magnitude lower than that of the human brain, 0.1—10 fJ per spike. Notably, this synaptic-like plasticity arises from the dynamic interplay between two cationic species (Na+ and K+), exemplifying an illustrative approach to leveraging ionic diversity for achieving neuromorphic functionality. The conductive state of the membrane is shown to be governed by the adsorption/desorption dynamics of Na+ trapped in the pores, while K+ ions permeate through the empty pores. We derived a set of illustrative analytical expressions associated with the apparent synaptic-like potentiation and memory volatility. Further experimentation is needed to demonstrate the predicted behaviors for various subnanoporous 2D membranes, as well as ionic permeants. It is our hope to stimulate further work aimed at a better understanding of the effects of ionic diversity, including at the system level, i.e., in the case of interconnected nanofluidic memristive elements.
Methods
@@
For the MD simulations of ion transport, we used a rectangular box with dimension $L_{X}=6.998$ nm $\times$ $L_{Y}=6.926$ nm $\times$ $L_{Z}=10$ nm, periodic in all directions. The subnanoporous hBN monolayer was placed in the $XY$-plane at $Z=L_{Z}/2$ with the pore array consisted of 16 triangular B3N vacancies. To prevent the membrane from drifting, its edge atoms were tethered to their initial positions by spring restraints. The system was filled with aqueous salt mixtures consisting of explicitly simulated K+, Na+, and Cl- ions. As the system was periodic in all directions, the time- and ensemble-averaged salt concentration on each side of the membrane was identical under zero bias. The membrane was simulated using parameters developed earlier Govind Rajan et al. (2018) within the OPLS-AA forcefield framework Jorgensen et al. (1996). The partial charges of the nitrogen atoms at the edges of B3N vacancies were set at $2/3$ of their bulk values (obtained earlier from quantum-chemical calculations Govind Rajan et al. (2018)), ensuring electrical neutrality of the membrane. Explicitly simulated water molecules were simulated using the TIP4P model Jorgensen et al. (1983). All of the non-bonded interactions were simulated using the OPLS-AA forcefield framework Jorgensen et al. (1996). A particle-particle–particle-mesh scheme was used to simulate the electrostatic interactions. A 1.2 nm cut-off radius was used for all short-range interactions, including short-range Coulomb interactions and the Lennard-Jones interactions. To initialize the conductive weight of the membrane to zero in all of our simulations, we placed Na+ ions inside the B3N vacancies. The system first underwent static energy minimization and then dynamics relaxation in the NPT ensemble at $T=300$ K and $P=1$ bar, using a 1 fs time step, with the Parrinello-Rahman barostat modifying simulation box dimensions only in the $Z$-direction. Relaxed systems underwent ion transport simulations under rectangular pulses of external electric field applied in the $Z$-direction, using a 2 fs time step. The corresponding pulse magnitudes were calculated as the electric field magnitude, multiplied by $L_{Z}$. To obtain the ion currents, we first calculated the cumulative ionic fluxes as a function of time $N(t)$. The flux data was recorded every 10 ps, corresponding to a 100 GHz sampling rate. The raw $N(t)$ data were then filtered using a Chebyshev low-pass digital filter with a 200 MHz cutoff frequency. Finally, time-derivatives of the filtered flux data were obtained using an 8th-order central difference method, yielding the time-dependent ionic current $I(t)=q\frac{dN_{f}(t)}{dt}$, where $N_{f}(t)$ is the filtered cumulative flux and $q$ is the ionic charge. Further details can be found in the supplementary materials of our previous work Noh and Smolyanitsky (2024a). The PMF data were obtained using the Weighted Histogram Analysis Method Hub et al. (2010) applied to a total of 60 umbrella samples of the ion’s position along the $Z$-coordinate incremented by 0.05 nm; each umbrella sample was obtained from a 20-ns-long simulation. All MD simulations were performed using GPU-accelerated GROMACS Abraham et al. (2015); Páll et al. (2020) and the molecular visualization tasks were carried using the OVITO software Stukowski (2009).
Acknowledgments
@@
A substantial portion of the simulations was conducted using the GPU-accelerated high-performance computing facilities at the National Institute of Standards and Technology. We thank Frances Allen for numerous illuminating discussions regarding fabrication of multivacancy arrays in 2D materials. Yechan Noh’s work on subdiffusive transport characterization was supported by National Science Foundation (NSF) under Award No. 2110924. Use of computational resources was supported by NSF under ACCESS program, Award No. NNT230006.